OpenStudy (anonymous):

how do i solve the equation by completing the square for x^2+6x=-8 with steps please

OpenStudy (irishboy123):

\(x^2+6x=-8\) first, can you write it as: \("something" \ = 0\)

OpenStudy (irishboy123):

where "something" has your various terms in it!

OpenStudy (anonymous):

wait what

OpenStudy (irishboy123):

can you make the RHS of your equation equal to zero

OpenStudy (anonymous):

what is that?

OpenStudy (irishboy123):

let's say i have : \(x^2+4x=−10\) i can then say that: \(x^2+4x \ + 10=−10 \ + \ 10\) \(x^2+4x + 10 = 0\) then i am good to go can tou do that with your equation?

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

wait what about the 4

OpenStudy (irishboy123):

what 4 ?

OpenStudy (anonymous):

wait 6x i meant

OpenStudy (irishboy123):

OK, we can complete the square without fiddling around with your equation so \(x^2+6x=-8\) look only at the L.H.S (left hand side): \(x^2+6x= (x + A)^2 - B\)

OpenStudy (anonymous):

wait is a and b the solution?

OpenStudy (irishboy123):

you will need A and B in the solution tracking back, you asked : "how do i solve the equation by completing the square for x^2+6x=-8 with steps please " that is a very reasonable question. do you know how to complete a square?

OpenStudy (anonymous):

no lol i got a 15 percent on my test and can not get it for the life of me

OpenStudy (irishboy123):

crumbs :-( what is \((x + 3)^2\) can you evaluate that?

OpenStudy (anonymous):

isit like (x+3) and (x-3)

OpenStudy (irishboy123):

crumbs almighty!! am i actually helping you or hindering you?!?! i do wonder !! \((x+3)^2 = (x+3)(x+3) = ....\)

OpenStudy (anonymous):

both sorta

OpenStudy (irishboy123):

ok love the directness!! am i now defunct?

OpenStudy (anonymous):

no sorry i was going somewhere

OpenStudy (irishboy123):

are we still going ???

OpenStudy (anonymous):

wait so i would add 8 to -8 and it would make it zero?